OpenStudy (anonymous):

Simplify the expression (2x − 9)(x + 6). 2x2 − 54 2x2 − 3x − 54 2x2 + 3x − 54 2x2 + 21x − 54

OpenStudy (demandandsupply):

Use the foil method of expansion

OpenStudy (anonymous):

If you use foil, you'll get C.

OpenStudy (y1066380):

Expend (2x − 9)(x + 6) =(2x^2 + 12x- 9x - 54) ....

OpenStudy (anonymous):

C is the ans (2x − 9)(x + 6) =2x^2 + 12x - 9x - 54 = 2x^2 + 3x -54